OpenStudy (anonymous):

Where would I begin doing 1/6k^2=1/3k^2-1/k

OpenStudy (anonymous):

Multiply both sides by the LCM

OpenStudy (misty1212):

\[\frac{1}{6k^2}=\frac{1}{3k^2}-\frac{1}{k}\]

OpenStudy (anonymous):

In this case: \[ \text{lcm}(6k^2,3k^2,k) = 6k^2 \]